Can you please show me the equation for f(x)=3x^2+18x+3
i need to know what the vertex is. I have been racking my brain all morning.

Answer by Alan3354(69443)

Can you please show me the equation for f(x)=3x^2+18x+3
That is the equation ??
-----------------
i need to know what the vertex is. I have been racking my brain all morning.
The x value for the vertex is -b/2a
at x = -18/6 = -3
------------------
f(-3) = 3*9 + 18*-3 + 3 = 27 - 54 + 3
f(-3) = -24 = y value at vertex
Vertex at (-3,-24)